OpenStudy (anonymous):

solve using any algebraic method x+3/4+y-1/3=1 and 2x-y=12 is the equation, please explain thanks

OpenStudy (anonymous):

I'm going to assume that you mean (x+3)/4+(y-1)/3=1 2x=12+y x=6+y/2 (6+y/2+3)/4+(y-1)/3=1 (9+y/2)/4+(y-1)/3=1 3(9+y/2)+4(y-1)=12 27+3y/2+4y-4=12 5.5y=-11 y=-2 2x=12-2 2x=10 x=5 (5,-2) is the point of intersection

OpenStudy (anonymous):

yes